OpenStudy (anonymous):

Verify that this equation is an identity. Cos(4u)=2cos^2(2u)-1

OpenStudy (anonymous):

\[\cos(4u)=2\cos^2(2u)-1\]

OpenStudy (anonymous):

Again using the double angle formula.. \[cos(4u) = cos^2(2u) - sin^2(2u)\] \[sin^2(\theta) + cos^2(\theta) = 1\] Just let \(\theta=2u\) and solve the bottom for \(sin^2\theta\) and plug into the top.

OpenStudy (anonymous):

I'm not done but \[\sin^2\theta=1-\cos^2\theta\]